OpenStudy (anonymous):

What is the value of x in the equation 0.7x – 1.4 = –3.5? A. –7 B. –3 C. 3 D. 7

OpenStudy (anonymous):

plug the numbers in the x tillu get 3.5

OpenStudy (anonymous):

0.7x=-3.5+1.4 0.7x=-2.1 x=-2.1/0.7 x=-3

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

Therefore b.-3 is your answer

OpenStudy (anonymous):

your welcome